In one room are 144 chairs. The number of chairs per row is two less than twice the number of rows. Find the number of rows and the number of chairs per row.
Answer by stanbon(75887)

In one room are 144 chairs. The number of chairs per row is two less than twice the number of rows. Find the number of rows and the number of chairs per row.
-------------
# of chairs = c
c = 144
Let # of rows be "r"::
Equation:
# of chairs per row = c/r
c/r = 2r-2
------
Solve for "r"::
144/r = 2r-2
----
2r^2 - 2r - 144 = 0
--------------
r^2 - r - 72 = 0
---
(r-9)(x+8) = 0
-----
r = 9 (# of rows)
---------------------
c/r = 144/9 = 16 (# of chairs per row)
Cheers,
Stan H.
-----------